OpenStudy (anonymous):

Does the triangle in 3 space with vertices (-1,2,3), (2,-2,0) and (3,1,-4) have an obtuse angle? How do I find the angles? How do I justify the answer?

Parth (parthkohli):

If\[c^2 > a^2 + b^2\]where \(c\) is the longest side and \(a,b\) are two others, then the triangle HAS an obtuse angle.

OpenStudy (anonymous):

So first I'll have to find all the distances?

Parth (parthkohli):

Yes!

OpenStudy (anonymous):

But this question is in my 3 Dimensional Space/Vectors chapter, I was wondering can there be any other method to solve the same question using vectors/matrix something like that?

Parth (parthkohli):

Finding the magnitude of a vector and finding the distance between two points is essentially the same thing.

Parth (parthkohli):

You can represent the vector which joins \(\left(-1,2,3\right)\) and \((2,-2,0)\) in component form as \(\langle 3,-4,3\rangle\) or \(3{\bf i} - 4{\bf j} + 3{\bf k}\). Find the magnitude of that.

Parth (parthkohli):

That thing would give you the length of one side. Similarly, represent the vector between \((2,-2,0)\) and \((3,1,-4)\) and then find the magnitude to get the length of the second side.

Parth (parthkohli):

Are you following me? @Zeerak

OpenStudy (anonymous):

Yes! Just worked the problem out!

OpenStudy (anonymous):

Greattt! Thanks for the thorough and prompt reply!

Parth (parthkohli):

Nice! So, what are the values of \(a,b,c\)?

OpenStudy (unklerhaukus):

why not just draw the triangle

Parth (parthkohli):

lol yeah, but that wouldn't qualify as "work" :-)

OpenStudy (unklerhaukus):

after you draw it , it is easy to see the sides can be found using Pythagorus

Parth (parthkohli):

@UnkleRhaukus Do you mean Law of Cosines? It's not a right triangle.

Parth (parthkohli):

In this case, we can't use the Law of Cosines either.

Parth (parthkohli):

We don't know the measure of all sides. :-|

Parth (parthkohli):

But a simple approach, as I said, is to check if \(c^2 > a^2 + b^2\)

OpenStudy (unklerhaukus):

oh, wait we are the 3space, (drawing is not going to be easy_

OpenStudy (anonymous):

a= rt34 b= rt66 c= rt26

Parth (parthkohli):

Remember that \(c\) should be \(\sqrt{66}\) since it's the LONGEST side.

OpenStudy (anonymous):

Oh yes, wrote b on that side

Parth (parthkohli):

Check if\[\left(\sqrt{66}\right)^2 > \left(\sqrt{34}\right)^2 + \left(\sqrt{26}\right)^2\]

OpenStudy (anonymous):

Yes 66 > 60

OpenStudy (anonymous):

It means it has an obtuse angle

Parth (parthkohli):

So it does have an obtuse triangle.\(\qquad \qquad \qquad \qquad \boxed{}\)

Parth (parthkohli):

Does the question ask for the angles as well?

OpenStudy (anonymous):

One last thing Parth, I tried to do it first by finding magnitude of (2,-2,0) and (-1,2,3) and then used costheta = u.v / ||u|| ||v|| but it gave me wrong answers

OpenStudy (unklerhaukus):

another way to find the angles is using the dot product

Parth (parthkohli):

I prefer the Law of Cosines but yeah.

OpenStudy (anonymous):

No, does not ask for the angles, but if I had to find the angles too, what should i do?

Parth (parthkohli):

You can use the dot product or the Law of Cosines... I think I like the Law of Cosines a bit too much lol :-P

OpenStudy (unklerhaukus):

yeah i would use this formula \[\cos\theta = \frac{u\cdot v}{ \|u\| \|v\|}\]

Parth (parthkohli):

Yeah, I'd use that formula too :-)

OpenStudy (anonymous):

Unkle, it gave me wrong answers

OpenStudy (unklerhaukus):

you have already worked out ||u|| and ||v||

OpenStudy (unklerhaukus):

i can't see your mistake if i can't see your working

Parth (parthkohli):

Optional:\[c^2 = a^2 + b^2 - 2ab\cos\gamma \]\[b^2 = a^2 + c^2 - 2ab\cos\beta\]\[a^2 = b^2 + c^2 - 2ab\cos \alpha\]So,\[66 =60 - 2\sqrt{34 \times 26}\cos\gamma \]\[34 = 92 - 2\sqrt{66 \times 26}\cos \beta \]\[26 = 100 - 2\sqrt{66 \times 34}\cos\alpha\]

Parth (parthkohli):

I'm being too annoying right now with that one. :-|

OpenStudy (anonymous):

That's absolutely fine! If I had to use he dot product method, without finding the lengths of the sides can that be done? What I did earlier: ||u|| = \[\sqrt{4+4}\] = rt8 = 2rt2 ||v|| = \[\sqrt{1+4=9}\] = rt14

OpenStudy (anonymous):

Because u = (2,-2,0) v = (-1,2,3)

Parth (parthkohli):

Well, the magnitude of the vectors are just the length of the sides in the triangle...

Parth (parthkohli):

Remember that \(u\) and \(v\) are vectors, not points.

OpenStudy (anonymous):

But Im not getting the same answers by two different methods? Why is that?

OpenStudy (anonymous):

Oh yes! u and v are vectors I just cant take the magnitude of two point and treat them as vectors!

OpenStudy (anonymous):

Oh Thank you Parth!

Parth (parthkohli):

You have to find the angle between two vectors. Do you remember the three vectors you got? I gave you one, \(\langle 3,-4,3\rangle \)

Parth (parthkohli):

You can't treat a point as a vector...

Parth (parthkohli):

How did you get the lengths of the three sides of the triangle?